# ChartOptimiser: Task-driven Optimisation of Chart Designs

### v1

@@ 3.2. Objective Function
The first metric $L_{w}$ is the White Space Ratio (WSR) (O’Donovan et al., 2014), also known as area usage (Wen et al., 2005). WSR is calculated as the ratio of pixels in $V_{\mathbf{x}}$ that are purely white (RGB value #FFFFFF) to the total number of pixels in $V_{\mathbf{x}}$. This metric is critical for quantifying the visual density of chart designs (Rosenholtz et al., 2007). A high WSR suggests the presence of visual clutter, while a low WSR implies inefficient space utilisation. The WSR distribution $\mathcal{N}(\mu,\sigma^{2})$, derived from human-crafted charts in the ChartQA dataset (Masry et al., 2022), is considered the golden standard. Any WSR that deviates significantly from this standard is penalised.
(2) $L_{w}(V_{\mathbf{x}})=\begin{cases}0&\text{ $WSR(V_{\mathbf{x}})\in(\mu-\sigma,\mu+\sigma)$}\\ -\left|WSR(V_{\mathbf{x}})-\mu\right|&\text{otherwise}\\ \end{cases}$
@@
The third metric is the Task Saliency ($L_{s}$), i.e. mean visual saliency in the task-related areas of interest (AOIs). It represents the average task-driven saliency value across the AOIs $task\_aoi_{i}(V_{x})$. A well-designed chart that supports task-solving should effectively direct viewers’ attention to these regions. Thus, a higher $L_{s}$ indicates more attention is drawn to the target region, suggesting better alignment with task objectives.
@@
The fourth metric, Text Legibility ($L_{t}$), evaluates the readability of text in visual designs (Shin et al., 2023). A design fails if the texts are not readable due to small font sizes, text overlap, or misalignment beyond the image region. We employ Optical Character Recognition (OCR) (Shi et al., 2016) to automatically detect text characters in the rendered charts to approximate human perception. A penalty is applied when OCR fails to detect the text, indicating poor legibility successfully. For all $M$ categorical and data labels, the OCR score equals 1 if a label $text_{m}$ is successfully detected and equals 0 if missing. To ensure a smoother objective function, we construct a $P$-level image pyramid, where $downsample_{p}(V_{\mathbf{x}})$ is a downsampled version of the original image by a factor of $p$.
@@
All parameters in the design space are constrained by at least one metric in the objective function (see Figure 3). The WSR controls the aspect ratio and bar width, preventing bar overlay or too far distances. The Colour Preference metric controls bar colours, making the colours in the chart more harmonious. The Task Saliency metric generally influences chart layouts (orientation, aspect ratio, bar width) and colour, improving colour harmony and bar distance. Finally, the Text Legibility constrains the font size of the category and data labels. For example, optimising for Text Legibility may influence the font size of categorical and data labels and the rotation of these labels, while optimising for WSR may influence the width and aspect ratio of the bars themselves. Optimising for both may influence all of these metrics jointly. The final objective function $L$ is
@@ 3.3. Bayesian Optimisation
Evaluating charts poses significant challenges as identifying the parameter hypercube that yields the optimal value of the objective function is time-consuming. The challenges are threefold. First, identifying a “golden standard” for a specific design parameter is hard since the inter-dependencies between parameters are inherently black-box. Second, despite using descriptive grammars like Vega-Lite, the design space $\mathcal{S}$ remains large, encompassing continuous chart design parameters such as bar width and colour. Third, each evaluation requires generating a visual representation of the chart, as all terms in the objective function depend on rendered images. For instance, saliency maps could only be predicted once images were rendered. Repeated image generation across iterations incurs substantial computational costs. These factors underscore the need for sample-efficient optimisation techniques to reduce the number of costly evaluations required to achieve an optimal layout.
Inspired by (Dudley et al., 2019) that used Bayesian optimisation to refine 2D interface design in a map searching use case, we built a BO pipeline to balance efficiency and effectiveness. Bayesian optimisation (Shahriari et al., 2015) is a popular approach to optimise costly functions across various applications, which typically employs a probabilistic surrogate model to approximate the black-box function and estimate uncertainty. An acquisition function then utilises these estimates to determine the next query point. The surrogate model is often a Gaussian process (GP) (Williams and Rasmussen, 2006), which uses a kernel or covariance function $k$ to encode the prior belief for the smoothness of the black-box function $f$.
@@
where $[K_{n}]_{ij}=k(\mathbf{x}_{i},\mathbf{x}_{j})$, $\mathbf{k}_{n}=[k(\mathbf{x},\mathbf{x}_{1}),\cdots,k(\mathbf{x},\mathbf{x}_{n})]^{\mathsf{T}}$ and $\sigma$ is the variance of the noise. $k(\mathbf{x},\mathbf{x^{\prime}})$ is the similarity prediction based on a type of kernel $k$. We use radial basis function kernel in the experiment. $\mathbf{k}_{n}$ is the vector containing all the similarity measures between input point x and all other existing data $x_{1},...,x_{n}$.
With the posterior prediction, we can compute the utility of a given point $\mathbf{x}$ via an acquisition function $\alpha(\mathbf{x})$. One of the most popular acquisition functions is the Expected Improvement (EI) (Mockus, 1974), which has the form:
@@
In summary, we use BO to find the optimal design parameters by treating the user’s perceptual behaviour as a black-box function. We use a GP (Williams and Rasmussen, 2006) to model the joint perceptual objective $L$ based on four perceptual terms described in section 3.2. The chart design parameters are optimised by maximising the acquisition function iteratively.
@@ 3.4. Implementation Details
Performing analytical tasks are common use cases when people read charts (Elzer et al., 2011; Polatsek et al., 2018). Good chart designs should let viewers easily solve analytical tasks for efficient information communication. In this work, four common analytical tasks were studied (see Figure 1):
  • Find Extremum (FE) (Amar et al., 2005). FE tasks ask for the label of a data point with an extreme value of an attribute. Example task: Which mode of transport has the fewest CO2 emissions? To solve this task, participants must find the shortest bar in the chart and read the corresponding categorical label Eurostar (international rail).
@@
  • Compare (CP) (Albers et al., 2014). CP tasks require participants to perform RV tasks to retrieve the value of given data points (usually two) and then make judgements about data properties. Example task: Does Domestic flight have more CO2 emission than Bus? To solve this task, participants should look up the values of domestic flights and buses and then judge whether domestic flights have more CO2 emissions.
Data Preparation
Implementing the proposed pipeline requires a vector-based chart dataset that supports customisable chart design parameters and includes task-based optimisation goals. We utilised the ChartQA dataset (Masry et al., 2022), a chart question-answering dataset comprising real-world charts and high-quality human-annotated questions. Each chart in ChartQA is associated with two human-annotated analytical tasks. We restricted our selection to single bar charts to simplify the design space. Since bar colours were changed in the optimisation pipeline, we filtered out visual questions about bar colours, such as "What is the value of the yellow bar?" The final selection comprised 300 bar charts (207 horizontal and 93 vertical). These charts were recreated using the Vega-Lite grammar (Satyanarayan et al., 2016) [^1] https://vega.github.io/vega-lite, which allowed for declarative control over chart properties via key-value pairs in JSON format. This reduces the design space to a reasonable scale, consistent with previous visualisation literature (Moritz et al., 2018; Wu et al., 2021). We created one horizontal and one vertical bar chart template with fixed design parameters in Vega-Lite and filled the JSON template with data values from the ChartQA annotations. In line with previous work (Wu et al., 2021), the chart height was fixed at 600 pixels, while the width was initialised to 600 pixels as well. Text elements were uniformly set to a font size of 15 pixels, bar widths were standardised to 40 pixels, and all bars were assigned the colour #949d48, a common hue in ChartQA. The orientation of each chart followed the original designs in ChartQA. The Vega-Lite compiler then rendered the JSON files into Portable Network Graphics (PNG) and Scalable Vector Graphics (SVG) formats.
Specification of Objective Functions
The implementation of the WAVE algorithm for colour preference was based on the Aalto Interface Metrics platform (Oulasvirta et al., 2018) [^2] https://github.com/aalto-ui/aim. Task-relevant regions were automatically detected from the rendered chart in SVG format. For the FE and RV tasks, one task-related region was identified. In contrast, two or more target regions were defined for the CDV and CP tasks. To align these regions with annotated questions or answers, we matched the "aria-label" attribute in the SVG with the corresponding elements. We utilised the boundingbox() function from the Python svg.path package to extract the bounding boxes. Task-driven saliency maps were generated using the VisSalFormer model (Wang et al., 2024c). Regions with zero saliency were excluded from the analysis by discarding pixels within bounding boxes that had a saliency value of 0. For text legibility assessments, we used Pytesseract [^3] https://pypi.org/project/pytesseract for OCR on a three-level image pyramid (at 1/8, 1/4, and 1/2 of the original image size).
Parameter Boundaries
The bounds for each parameter were carefully selected to define the limits of the hypercube. We set the aspect ratio between 0.33 and 3, varying the chart width from 200 to 1,800 pixels. Font sizes were restricted to 10-36 pixels, bar width to 20 - 180 pixels, and label rotation to {0, -45, -90}. We permitted the full range of the HSV colour space for bar colours.
Training Details
Weights of the objective function were set to $w_{s}=4$, $w_{c}=1$, $w_{w}=3$, $w_{t}=2$. The candidate list was constructed by sampling from the parameter hypercube using a Sobol sequence. We sampled 50 candidates with ChartOptimiser for every chart-task pair, which took about 5-7 seconds per iteration. While more candidates provide better search resolution, we considered 50 candidates sufficient to demonstrate the strength of ChartOptimiser. All experiments were conducted on a Desktop PC with an AMD Ryzen 7 3700K 8-Core Processor and NVIDIA GeForce RTX 2060 Super GPU with 8GB VRAM.
@@ 4. Evaluation
To gain further insights, we conducted a user study to let participants qualitatively rate chart designs generated from our method against human-designed charts and three strong baseline approaches. Then, we further analysed the optimised parameters from ChartOptimiser.
@@ 4.1. Study Design
We randomly sampled 12 charts from the ChartQA dataset, corresponding to three tasks from each of the four task types introduced in Section 3: Find Extremum (FE), Retrieve Value (RV), Compute Derived Value (CDV), and Compare (CP). RV and FE were considered simple tasks, as they require observing only a single data point. In contrast, CP and CDV were considered more complex, necessitating observing multiple data points and performing calculations (Elzer et al., 2011). In the study, participants were introduced to the tasks before starting. We presented participants with five chart designs in each of the 12 study tasks (see Figure 4). Study participants were asked to evaluate the five chart designs from three aspects: 1) chart aesthetics, 2) chart clarity, 3) and task-solving ease. Each aspect was evaluated on a 7-point Likert scale, where 1 is the strongest negative, and 7 is the strongest positive. After completing ratings for each task, participants are asked to leave comments to briefly explain their ratings.
@@
  • Vega-Lite default designs (VegaLite). These are the recreated charts from ChartQA data tables, using the Vega-Lite library (Satyanarayan et al., 2016) with fixed chart design parameters (see Section 4.2).
  • GPT-4o optimised charts (GPT-4o). These are the responses from GPT-4o [^4] https://platform.openai.com/docs/models/gpt-4o (Achiam et al., 2023) with the prompt “{\$VEGA_JSON}”, “optimise the above Vega-lite JSON with the task: {\$TASK}”. The {\$VEGA_JSON} is the JSON file of the VegaLite baseline, while the {\$TASK} is either FE, RV, CP, or CDV task from the ChartQA dataset. We noticed that GPT-4o filters data in some CP and CDV tasks, so we added the prompt “please add the filtered data back” to prevent editing the data source.
@@
Figure 4. From left to right: Sample bar charts from the ChartQA (Masry et al., 2022) dataset as well as corresponding bar charts optimised using ChartOptimiser, Vega-Lite default (Satyanarayan et al., 2016), LQ2 (Wu et al., 2021), and GPT-4o (Achiam et al., 2023).
[…]
Sample bar charts from the ChartQA dataset, optimised bar charts from ChartOptimiser, Vega-Lite default, LQ2, and GPT-4o.
@@ 4.2. Study Results
Figure 5 illustrates the participants’ ratings for visual aesthetics, chart clarity, and task-solving ease for all chart designs. Overall, ChartOptimiser ranks first in clarity ($M=5.6,SD=1.4$) and task-solving ease ($M=6.1,SD=1.2$), with the GPT-4 design coming in second. The human-designed chart outperforms others in aesthetics ($M=5.4,SD=1.5$), with our ChartOptimiser ranking second ($M=4.9,SD=1.5$). The results indicate that human designers consider more details to ensure the charts are pleasing, while automatic approaches can improve chart clarity and task-solving ease. The normality of the ratings was confirmed via the Shapiro-Wilk Normality test. A one-way ANOVA revealed a statistically significant difference in ratings between these methods in aesthetic ($F(4,1795)=39.2,p<0.01$) and clarity ($F(4,1795)=3.2,p<0.01$). The task-solving ease was insignificant based on the one-way ANOVA analysis ($F(4,1795)=1.8,p=0.13$). No significant differences were found in the post-hoc Tukey’s HSD test between methods regarding task-solving ease (p > 0.05). However, in terms of aesthetics, human designs were significantly better than other methods (p < 0.05). Additionally, ChartOptimiser was significantly better than LQ2 (p < 0.05) and slightly better than GPT-4o and VegaLite (p > 0.05). In clarity, ChartOptimiser was significantly better than human designs and LQ2 (p < 0.05), but also slightly higher than GPT-4o and VegaLite (p > 0.05). This result indicates that the method has the potential to design charts suitable for tasks in terms of both aesthetics and clarity.
@@
Next, we examined the ratings for each task (see Figure 6): 1) CDV: ChartOptimiser outperformed others in clarity and task-solving ease, ranked second in aesthetics, following human-designed charts. 2) CP: ChartOptimiser ranked first in aesthetics, surpassing human designs that ranked second. This is the only task for which human designs did not rank first. Our method ranked in the top two for clarity and task-solving, with no significance found against LQ2, which ranked first (p > 0.05). 3) FE: All designs received high task-solving ratings, suggesting this task’s easiness. VegaLite ranked first in clarity, as the simplest design can show extreme values clearly. Human design ranked first in aesthetics. 4) RV: GPT-4o designs performed well in the simple retrieve value task, with ChartOptimiser in second place with no significance. Our proposed approach performed best in complex tasks (CDV and CP). It also demonstrated its superiority in simple tasks, although comparable to other approaches.

### v2

@@ 3.2. Objective Function
The first metric $L_{w}$ is the White Space Ratio (WSR) (O’Donovan et al., 2014), also known as area usage (Wen et al., 2005). WSR is calculated as the ratio of pixels in $V_{\mathbf{x}}$ that are purely white (RGB value #FFFFFF) to the total number of pixels in $V_{\mathbf{x}}$. This metric is critical for quantifying the visual density of chart designs (Rosenholtz et al., 2007). A high WSR suggests the presence of visual clutter, while a low WSR implies inefficient space utilisation. The WSR distribution $\mathcal{N}(\mu,\sigma^{2})$, derived from human-crafted charts in the ChartQA dataset (Masry et al., 2022), is considered the golden standard. Any WSR that deviates significantly from this standard is penalised for visually cluttered designs or leaving too much white space.
[…]
@@
The third metric is the Task Saliency ($L_{s}$), i.e. mean visual saliency in the task-related areas of interest (AOIs). This is inspired by previous works optimising graphical designs with visual importance maps (O’Donovan et al., 2014; Bylinskii et al., 2017; Fosco et al., 2020). This term represents the average task-driven saliency value across the AOIs $task\_aoi_{i}(V_{x})$. A well-designed chart that supports task-solving should effectively direct viewers’ attention to these regions. Thus, a higher $L_{s}$ indicates more attention is drawn to the target region, suggesting better alignment with task objectives.
@@
The fourth metric, Text Legibility ($L_{t}$), evaluates the readability of text in visual designs (Shin et al., 2023). A design fails if the text is not readable due to small font sizes, text overlap, or misalignment that extends beyond the image region. We employ Optical Character Recognition (OCR) (Shi et al., 2016) to automatically detect text characters in the rendered charts to approximate human perception. A penalty is applied when OCR fails to detect the text, indicating poor legibility. For all $M$ categorical and data labels, the OCR score equals 1 if a label $text_{m}$ is successfully detected and equals 0 if missing. To ensure a smoother objective function, we construct a $P$-level image pyramid, where $downsample_{p}(V_{\mathbf{x}})$ is a downsampled version of the original image by a factor of $p$.
@@
All parameters in the design space are constrained by at least one metric in the objective function (see Figure 3). The WSR controls the aspect ratio and bar width, preventing bar overlay or too far distances. The Colour Preference metric controls bar colours, making the colours in the chart more harmonious. The Task Saliency metric generally influences chart layouts (orientation, aspect ratio, bar width) and colour, improving colour harmony and bar distance. Finally, the Text Legibility constrains the font size of the category and data labels. For example, optimising for Text Legibility may influence the font size of categorical and data labels and the rotation of these labels, while optimising for WSR may influence the width and aspect ratio of the bars themselves. Optimising for both may influence all of these metrics jointly.
The final objective function $L$ is
@@ 3.3. Bayesian Optimisation
Evaluating charts poses significant challenges as identifying the parameter hypercube that yields the optimal value of the objective function is time-consuming. The challenges are threefold. First, identifying a “golden standard” for a specific design parameter is challenging since the interdependencies between parameters are inherently complex. Second, despite using descriptive grammars like VegaLite, the design space $\mathcal{S}$ remains large, encompassing continuous chart design parameters such as bar width and colour. Third, each evaluation requires generating a visual representation of the chart, as all terms in the objective function depend on rendered images. For instance, saliency maps could only be predicted once images were rendered. Repeated image generation across iterations incurs substantial computational costs. These factors underscore the need for sample-efficient optimisation techniques to reduce the number of costly evaluations required to achieve an optimal layout.
Inspired by (Dudley et al., 2019) that used Bayesian optimisation to refine 2D interface design in a map searching use case, we built a BO pipeline to balance efficiency and effectiveness. Bayesian optimisation (Shahriari et al., 2015) is a popular approach to optimise costly functions across various applications, which typically employs a probabilistic surrogate model to approximate the black-box function and estimate uncertainty. An acquisition function then utilises these estimates to determine the next query point. The non-parametric and sample-efficient nature of Gaussian Process fits our design optimisation problem well. Therefore, we use a Gaussian process (GP) (Williams and Rasmussen, 2006) as the surrogate model, which uses a kernel or covariance function $k$ to encode the prior belief for the smoothness of the black-box function $f$.
@@
where $K_{n}]_{ij}=k(\mathbf{x}_{i},\mathbf{x}_{j})$, $\mathbf{k}_{n}=[k(\mathbf{x},\mathbf{x}_{1}),\cdots,k(\mathbf{x},\mathbf{x}_{n})]^{\mathsf{T}}$ and $\sigma$ is the variance of the noise. $k(\mathbf{x},\mathbf{x^{\prime}})$ is the similarity prediction based on a type of kernel $k$. We use radial basis function kernel in the experiment. $\mathbf{k}_{n}$ is the vector containing all the similarity measures between input point x and all other existing data $x_{1},...,x_{n}$.
With the posterior prediction, we can compute the utility of a given point $\mathbf{x}$ via an acquisition function $\alpha(\mathbf{x})$. One of the most popular acquisition functions is the Expected Improvement (EI) (Mockus, 1974):
@@
In summary, we use BO to find the optimal design parameters by treating the user’s perceptual behaviour as a black-box function. We use a Gaussian Process (Williams and Rasmussen, 2006) to model the joint perceptual objective $L$ based on four perceptual terms described in section 3.2. The chart design parameters are optimised by maximising the acquisition function iteratively.
@@ 3.4. Implementation Details
Specification of Objective Functions
For white space ratio, the normal distribution has $\mu=0.496,\sigma=0.263$. The implementation of the WAVE algorithm for colour preference was based on the Aalto Interface Metrics platform (Oulasvirta et al., 2018) [^1] https://github.com/aalto-ui/aim. For task saliency, task-relevant regions were automatically detected from the rendered chart in SVG format. For the FE and RV tasks, one task-related region was identified. In contrast, two or more target regions were defined for the CDV and CP tasks. To align these regions with annotated questions or answers, we matched the “aria-label” attribute in the SVG with the corresponding elements. We utilised the boundingbox() function from svg.path [^2] https://pypi.org/project/svg.path/ to extract the task-related bounding boxes. The saliency maps were generated using the VisSalFormer model (Wang et al., 2024d). Regions with zero saliency were excluded from the analysis by discarding pixels within bounding boxes that had a saliency value of 0. For text legibility assessments, we applied Optical Character Recognition (OCR), using PyTesseract [^3] https://pypi.org/project/pytesseract on a three-level image pyramid (at 1/8, 1/4, and 1/2 of the original image size).
Parameter Boundaries
The bounds for each parameter were carefully selected to define the limits of the hypercube. We set the aspect ratio between 0.33 and 3, varying the chart width from 200 to 1,800 pixels. Font sizes were restricted to 10-36 pixels, bar width to 20-180 pixels, and label rotation to {0, -45, -90}. We permitted the full range of the HSV space for bar colours.
Training Details
The BO pipeline is built with the BoTorch [^4] https://botorch.org/ and Adaptive Experimentation Platform [^5] https://ax.dev/. Weights of the objective function were set to $w_{w}=3$, $w_{c}=1$, $w_{t}=2$, $w_{s}=4$. The candidate list was constructed by sampling from the parameter hypercube using a Sobol sequence, with 16 trials generated and a minimum of 5 trials completed in each step. We sampled 50 candidates with ChartOptimiser for every chart-task pair, which took about 3-6 seconds per step. While more candidates provide better search resolution, we considered 50 candidates sufficient to demonstrate the strength of ChartOptimiser. All experiments were conducted on a Desktop PC with an AMD Ryzen 7 3700K 8-Core Processor and NVIDIA GeForce RTX 2060 Super GPU with 8GB VRAM.
4. Dataset
Data Preparation
Implementing the proposed pipeline requires a vector-based chart dataset that supports customisable chart design parameters and includes task-based optimisation goals. We utilised the ChartQA dataset (Masry et al., 2022), a chart question-answering dataset comprising real-world charts and high-quality human-annotated questions. Each chart in ChartQA is associated with two human-annotated analytical tasks. We restricted our selection to single- and multi-column bar charts to simplify the design space. Since bar colours were changed in the optimisation pipeline, we filtered out visual questions about bar colours, such as “What is the value of the yellow bar?”. The final selection comprised 300 single-column bar charts (207 horizontal and 93 vertical) and 258 multi-column bar charts (all horizontal). These charts were recreated using the VegaLite grammar (Satyanarayan et al., 2016) [^6] https://vega.github.io/vega-lite, which allowed for declarative control over chart properties via key-value pairs in JSON format. This reduces the design space to a reasonable scale, consistent with previous automatic design literature (Moritz et al., 2018; Wu et al., 2021). We created templates for single-column and multi-column bar charts with fixed design parameters in VegaLite and filled the JSON template with data values from the ChartQA annotations. In line with previous work (Wu et al., 2021), the chart height was fixed at 600 pixels, while the width was initialised to 600 pixels. Bar widths were all initiated to 40 pixels, and data labels and axis labels were uniformly set to a font size of 24 and 17 pixels, respectively. For single-column bars, all bars were assigned the colour #949d48, a common hue in ChartQA. For multi-column bars, we used the default colour palette in VegaLite. The orientation of each chart followed the original designs in ChartQA. The VegaLite compiler then rendered the JSON files into Portable Network Graphics (PNG) and Scalable Vector Graphics (SVG) formats, which were generated by Vega-Altair  [^7] https://altair-viz.github.io/.
@@ 4. Dataset
Performing analytical tasks is a common use case when people read charts (Elzer et al., 2011; Polatsek et al., 2018). Effective chart designs should enable viewers to easily perform analytical tasks, facilitating efficient communication of information. In this work, four common analytical tasks were studied (see Figure 1):
  • Find Extremum (FE) (Amar et al., 2005). FE tasks require labelling a data point with an extreme value of an attribute. Example task: Which mode of transport has the fewest CO2 emissions? To solve this task, participants must find the shortest bar in the chart and read the corresponding categorical label Eurostar (international rail).
@@
  • Compare (CP) (Albers et al., 2014). CP tasks require participants to perform RV tasks to retrieve the value of given data points (usually two) and then make judgements about data properties. Example task: Does Domestic flight have more CO2 emission than Bus? To solve this task, participants should look up the values for domestic flights and buses and then determine whether domestic flights have higher CO2 emissions.
@@ 5. Evaluation
To gain further insights, we conducted a user study to let participants qualitatively rate chart designs generated from our method against human-designed charts and three strong baseline approaches. Then, we further analysed the optimised parameters from ChartOptimiser, and showcased how ChartOptimiser could be applied to other chart types, including multi-column bar and circular charts.
@@ 5.1. Study Design
We randomly sampled 12 charts from the ChartQA dataset, corresponding to three tasks from each of the four task types introduced in Section 3: Find Extremum (FE), Retrieve Value (RV), Compute Derived Value (CDV), and Compare (CP). RV and FE were considered simple tasks, as they require observing only a single data point. In contrast, CP and CDV were considered more complex, necessitating observing multiple data points and performing calculations (Elzer et al., 2011). In the study, participants were introduced to the tasks before starting. We presented participants with five chart designs in each of the 12 study tasks (see Figure 4). Study participants were asked to evaluate the five chart designs from three aspects: 1) chart aesthetics, 2) chart clarity, 3) and task-solving ease. Each aspect was evaluated on a 7-point Likert scale, where 1 is the strongest negative, and 7 is the strongest positive. After completing ratings for each task, participants are asked to provide brief comments explaining their ratings.
@@
  • VegaLite default designs (VegaLite). These are the recreated charts from ChartQA data tables, using the VegaLite library (Satyanarayan et al., 2016) with fixed chart design parameters (see Section 4.2).
  • GPT-4o optimised charts (GPT-4o). These are the responses from the API of GPT-4o [^8] https://platform.openai.com/docs/models/gpt-4o (Achiam et al., 2023) with the prompts “You are an expert in data visualisation. Given the following VegaLite chart specification, suggest improvements to enhance chart clarity and visual aesthetics. VegaLite Specification: {\$VEGA_JSON}”, and “also optimise the above Vegalite JSON with the task: {\$TASK}, respond only with the new VegaLite Specification in JSON format”. The {\$VEGA_JSON} is the JSON file of the VegaLite baseline, while the {\$TASK} is either FE, RV, CP, or CDV task from the ChartQA dataset. We noticed that GPT-4o filters data in some CP and CDV tasks, so we added the prompt “do not change the data, add the filtered data back” to prevent editing the data source.
@@
Figure 4. From left to right: Sample bar charts from the ChartQA (Masry et al., 2022) dataset and corresponding bar charts optimised using ChartOptimiser, GPT-4o, LQ2 (Wu et al., 2021), as well as VegaLite default.
Sample bar charts from the ChartQA dataset, optimised bar charts from ChartOptimiser, LQ2, and GPT-4o, as well as VegaLite default.
@@ 5.2. Study Results
Figure 5 illustrates the participants’ ratings for visual aesthetics, chart clarity, and task-solving ease for all chart designs. Overall, ChartOptimiser ranks first in clarity ($M=5.6,SD=1.4$) and task-solving ease ($M=6.1,SD=1.2$), with the GPT-4 design coming in second. The human-designed chart outperforms others in aesthetics ($M=5.4,SD=1.5$), with our ChartOptimiser ranking second ($M=4.9,SD=1.5$). The results indicate that human designers consider more details to ensure the charts are pleasing, while automatic approaches can improve chart clarity and task-solving ease. The normality of the ratings was confirmed via the Shapiro-Wilk Normality test. A one-way ANOVA revealed a statistically significant difference in ratings between these methods in aesthetic ($F(4,1795)=39.2,p<0.01$) and clarity ($F(4,1795)=3.2,p<0.01$). The task-solving ease was insignificant based on the one-way ANOVA analysis ($F(4,1795)=1.8,p=0.13$). No significant differences were found in the post-hoc Tukey’s HSD test between methods regarding task-solving ease ($p>0.05$). However, in terms of aesthetics, human designs were significantly better than other methods ($p<0.05$). Additionally, ChartOptimiser was significantly better than LQ2 ($p<0.05$) and slightly better than GPT-4o and VegaLite ($p>0.05$). In clarity, ChartOptimiser was significantly better than human designs and LQ2 ($p<0.05$), but also slightly higher than GPT-4o and VegaLite ($p>0.05$). This result suggests that the method has the potential to design charts that are both aesthetically pleasing and clear for tasks.
@@
Next, we examined the ratings for each task (see Figure 6): 1) CDV: ChartOptimiser outperformed others in clarity and task-solving ease, ranked second in aesthetics, following human-designed charts. 2) CP: ChartOptimiser ranked first in aesthetics, surpassing human designs that ranked second. This is the only task for which human designs did not rank first. Our method ranked in the top two for clarity and task-solving, with no significance found against LQ2, which ranked first ($p>0.05$). 3) FE: All designs received high task-solving ratings, suggesting this task’s easiness. VegaLite ranked first in clarity, as its simplest design can clearly display extreme values. Human design ranked first in aesthetics. 4) RV: GPT-4o designs performed well in the simple retrieve value task, with ChartOptimiser in second place with no significance. Our proposed approach performed best in complex tasks (CDV and CP). It also demonstrated its superiority in simple tasks, although comparable to other approaches.